A box rests on a frozen pond, which serves as a frictionless

horizontal surface. If a fisherman applies a horizontal force with magnitude 48.0 N to the box and produces an acceleration of magnitude 3.00 m/s2, what is the mass of the box?


1
Expert's answer
2021-12-12T16:42:50-0500

Solution.

"F=48.0 N;"

"a=3.00 m\/s^2;"

"m-?;"

"a=\dfrac{F}{m}\implies m=\dfrac{F}{a};"

"m=\dfrac{48.0N}{3.00 m/s^2}=16.0 kg;"

Answer: "m=16.0 kg."